Use the quadratic formula to solve the equation. I forgot how to do this. Can someone please help me
x^2 - x = -1

Answer by stanbon(75887)

Use the quadratic formula to solve the equation. I forgot how to do this. Can someone please help me
x^2 - x = -1
Rearrange:
x^2 -x + 1 = 0
a = 1 ; b = -1 ; c = 1
---
Formula: x = [-b+-sqrt(b^2-4ac)]/(2a)
---
Your Problem:
x = [1 +- sqrt(1 -4*1*1)]/(2*1)
----
x [1 +- sqrt(-3)]/2
---
x = (1/2) + (i/2)sqrt(3) or x = (1/2) - (i/2)sqrt(3)
========================================================
cheers,
Stan H.